# PEA-Diffusion: Parameter-Efficient Adapter with Knowledge Distillation in non-English Text-to-Image Generation

### v1

@@ 4.1 Data Preparation
We download the Laion2B-multi [30] dataset along with English translations for image descriptions. Based on the open-source CLIP models available for different languages: Chinese CLIP, Russian CLIP, Japanese CLIP, Korean CLIP, and Italian CLIP, we determine to conduct experiments on the above five languages. For Chinese, we also collect Wukong [7] data with Chinese cultural image-text pairs and Chinese-English parallel corpus translations of Laion2B-en. For more data, model and code details, please refer to Appendix 6.
@@ 4.2 Implementation Details and Evaluation
Implementation details. We conduct all experiments on Stable Diffusion (SD)- [26],and Stable Diffusion XL (SDXL)- [22] like backbones. Take SDXL base model as an example. PEA-Diffusion consists of VAE, UNet, CLIP text encoder, and the adapter module, comprising 3 billion parameters, out of which a mere 6 million parameters ($\approx$ 0.2%) are trainable. The entire model is trained on 64 A100 GPUs for 70000 steps with a batch size of 5 per GPU.
@@
Evaluation metrics. For the general evaluation data MG, the widely used FID metric can only measure the feature distance between generated images and specific images. It often fails to provide accurate assessment for high-quality generated images, as emphasized and experimentally validated by SDXL [22] and other studies. Therefore, in order to evaluate the quality of generated images more accurately, we introduce two human preference models as our evaluation metrics: ImageReward [32] and PickScore [14]. Additionally, we continue to utilize OpenCLIP ViT-bigG to measure the similarity between images and text. As for the language-specific evaluation data MC, constructed for each language, to ensure a fair comparison of content relevance, we only consider the CLIP model trained on the specific language to calculate the correlation between generated images and their corresponding text. Since the general CLIP model can only process non-English data after translation into English, this process can introduce significant performance degradation. Furthermore, our paper’s focus is precisely on addressing the errors caused by translating non-English prompts into English to generate images. Therefore, other metrics that involve translation processes, such as the two human preference-based models mentioned above, will no longer be valid for evaluation.
@@
  • Fine-tuning with LoRA [12] and $W_{K},W_{V}$ (key and value projection matrices): As low-cost experiments for comparison, we also finetuning the model to transfer language capabilities.
@@ 4.3 Experimental Results
Table 1: CLIPScore with different specific languages. For Chinese, we evaluated two general evaluation metrics with MG data. $\dagger$ Indicates that it was tested based on its open source model.
@@
Altdiffusion m9 [35$\dagger$ 0.2100 -0.6966 0.1202 0.3715 0.3177 0.1906 0.2461
Altdiffusion m18 [35$\dagger$ 0.2172 0.3316 0.1648 0.4730 0.3502 0.1917 0.3008
GlueGen [23$\dagger$ 0.1700 -0.5187 0.1224 0.4184 0.3065 0.1769 0.2506
@@
Generation analysis. We validate our approach based on two different T2I base models. Firstly, to ensure a fair comparison with GlueGen and AltDiffusion, we train PEA-Diffusion based on the SD1.5 model. Secondly, to validate our low-cost experiment approach, we train a larger model based on SDXL. The compared methods are LoRA finetuning and $W_{K},W_{V}$ finetuning, both of which are parameter-efficient finetuning strategies. For the second approach, as proposed in [15], when updating the mapping from given text to image distribution, only updating $W_{K}^{(i)},W_{V}^{(i)}$ in each cross-attention block $i$ is sufficient since text features are the only input to the key and value projection matrix. Furthermore, to validate the wide application of the proposed approach with different text encoders, we train our model with a multi-language T5 encoder and a multi-language CLIP encoder. This type of training only requires English text data, which makes it even simpler and more efficient. Fig. 2 illustrates comparisons among translation and our methods with different training strategies. It can be observed that our PEA-Diffusion can achieve Chinese cultural-related image generation, but translation totally fails in these cases.
@@
As shown in the comprehensive results of Table 1, when compared with the translation method, our PEA-Diffusion outperforms it for almost all languages (except Japanese) in terms of CLIPScore for MC datasets. It is obvious that the translation method cannot capture the essence of different languages’ cultures, leading to a semantic gap between the original prompts and the translated English prompts. However, our approach can use additional cultural-related image-text pairs to train the adapter, which can align the language understanding ability of the text encoder with the generation ability of the UNet. Even though the UNet is frozen during the whole training process, we argue that our PEA-Diffusion can still stimulate the potential generation ability for cultural-related images of the original UNet. Since SD1.5 and SDXL are all trained on billions of image-text pairs that contain wide image domains, our approach can successfully awaken their generative abilities in different cultural contexts. In terms of ImageReward and PickScore for the MG dataset, we must admit that PEA-Diffusion is slightly inferior to translation methods. This is because using other text encoders (e.g., Chinese CLIP and Multilingual CLIP) cannot ensure complete alignment between the text embedding and the UNet feature spaces.
@@
Even though Altdiffusion retrains the multilingual CLIP text encoder and UNet on billions of image-text pairs, it still performs worse than our approach in terms of almost all metrics for all languages. GlueGen is the most related work on our concerned topic. The statistics presented in Table 1 reveal that both checkpoints provided by the authors or re-implemented by ourselves are significantly inferior to our PEA-Diffusion, especially for the ImageReward and PickScore metrics, resulting in very low generation quality since only aligning distributions of text embedding for different languages is very hard, even with its sophisticatedly designed GlenNet. In addition, it cannot generate cultural-related images. Therefore, it’s worth mentioning that our core contribution is that we align the model’s generation ability through UNet distillation only by a MLP-like adapter between different languages. We don’t forcibly align the text embedding spaces of different languages.
Further, we also conduct several parameter-efficient finetuning approaches for comparisons. Both LoRA and $W_{k}$ and $W_{v}$ finetuning methods perform poorly in terms of ImageReward, PickScore and CLIPScore metrics. This is because the aforementioned finetuning methods heavily rely on high-quality image-text datasets, whereas our PEA-Diffusion requires only some parallel multilingual texts and images without strict quality requirements.
@@
Analysis on different languages. The statistics presented in Table 1 show that the CLIPScores of Russian, Italian, and Korean are on par with or surpass other baseline methods significantly. Only for Japanese, the CLIPScore is slightly lower than that of translation and AltDiffusion. We attribute this to the weak encoding capability of the CLIP model for Japanese that cannot be aligned with the SD model through the adaoter layer. Besides, the training samples for Korean are only 1.6M, further demonstrating that our method can yield satisfactory language transfer without large-scale training datasets. More visualization examples for cultural-related generation results in different languages are illustrated in Appendix 7.3.
@@
Analysis of training cost Our PEA-Diffusion freezes the CLIP text encoder, the UNet, and VAE during the whole training process; only the MLP mapping adapter is trained, accounting for merely 6M training parameters, which is 0.2% of total parameters for SDXL-based models. We employed 64 A100 GPUs to train 70000 steps, resulting in 1600 GPU hours. As a comparison, AltDiffusion [35] trained the model for three stages with nearly 30000 GPU hours, which is $\approx 20$ times our training cost. In addition, our method achieves higher performance in terms of almost all metrics and all five experimental languages. We also re-implemented and trained GlueGen [23] with the same dataset constructed by ourselves. It indeed achieves very low training costs with only 1 A100 GPU for 24 hours, but results in very poor performance no matter for image quality (-0.5 ImageReward score) or language transferring ability. Finally, we also conduct other parameter-efficient finetuning strategies (e.g., LoRA and tuning only $W_{k}$ and $W_{v}$ matrices of attention layers) and achieve much lower performance and higher training costs, which demonstrated the effectiveness and efficiency of our PEA-Diffusion.
@@
Finetuning -1.3154 0.1854 0.3278 0.1901
Finetuning w/ TKD -0.9875 0.1997 0.3506 0.2287
Finetuning w/ ULKD -1.0447 0.1957 0.3331 0.1968
Finetuning w/ UFKD 0.9556 0.2174 0.4085 0.2557
Finetuning w/ UTKD 0.8490 0.2113 0.4011 0.2461
Ours(Finetuning w/ UKD) 1.0245 0.2244 0.4224 0.2610
4.4 Ablation Studies
Since Chinese data has certain advantages in terms of data quality and quantity, to eliminate errors caused by data, our ablation experiments are all based on Chinese data. The experiments include five parts: First, we conduct a baseline experiment where we only perform standard model fine-tuning without incorporating any KD strategies. Secondly, inspired by GlueGen [23], we employ text encoder KD to explore the impact of distillation at different positions in the SD framework. Next, we perform distillation based on the UNet model, including feature distillation between different UNet layers and logit distillation for the final output. Finally, we combine all of the distillation positions described above, including both text encoder and UNet KDs.
Effectiveness of KD. Firstly, KD plays an essential role in directing the model towards the correct learning path to quickly converge, and it is an indispensable method for PEA-Diffusion. As shown in the first line of Table 2, the pure finetuning approach results in significantly poor performance in terms of image quality (ImageReward and PickScore) and image-text matching degree (CLIPScores).
@@ 4.4 Ablation Studies
Effectiveness of UNet Feature KD (UFKD). Our KD method consists of two parts: the first part involves distilling the logit results generated by the UNet (ULKD), which is a relatively weak form of distillation. Therefore, we introduce a distillation of intermediate layer features in the UNet, which serves as the primary knowledge guidance mechanism. As shown in the third and fourth lines of Table 2, we can observe that logit distillation can only improve marginally when compared with direct finetuning whereas feature distillation boosts the performance remarkably. We attribute this phenomenon to the generation ability of the large amount of UNet parameters so that intermediate feature supervision is essential for knowledge transfer.
@@
Effectiveness of UNet KD. Furthermore, different distillation positions have different effects. Our initial focus is on aligning the feature spaces of different text encoders, so a natural idea is to distill only the feature information of the teacher model’s text encoder (TKD). However, as the results show in the second line of Table 2, although the overall performance exceeds the pure finetuning method, it is still far from our best-performing approach. This illustrates that simply aligning the text encoder is insufficient and ineffective. This statistic also matches the experimental results of GlueGen in Table 1: simply aligning text embedding spaces cannot obtain satisfactory image generation results. Therefore, this further proves the necessity of distilling the features of the UNet itself.
@@
Finally, we perform KD on both the UNet and text encoder (UTKD). However, the fifth line of Table 2 shows that its effect is worse than performing only UNet distillation. We suspect that the addition of text encoder KD actually weakens the learning of the SD core UNet structure. The main knowledge ability for generating images is primarily influenced by the UNet model, which is accountable for the adapter alignment feature in PEA-Diffusion.
Figure 3: Downstream task experiments of our proposed PEA. By only adding the trained adapter between the Chinese CLIP text encoder and the UNet, it can achieve language transfer for finetuning, LoRA, ControlNet, Inpainting, LCM and compressed models.
@@ 4.5 Adaptability Experiment of English Downstream Model
In this section, we would like to emphasize the plug-and-play feature of our PEA module. Currently, a wide range of community resources, such as fine-tuned models, LoRA, Dreambooth, ControlNet, compressed SD models, and accelerated sampling SD models, are derived from the English SD models, so that the English community is much maturer than other languages. Therefore, one key challenge we must overcome is bridging the compatibility gap between the new model and the English SD community after language transfer. Since our approach uses feature-level and logit-level KD to train only the PEA module while keeping the core UNet parameters unchanged, it naturally maintains compatibility with the English SD community and can seamlessly integrate with various English SD community plugins. Interestingly, the PEA module remains plug-and-play for models that have undergone changes in base SD UNet parameters in contrast to PEA-Diffusion but whose overall structure remains similar. This includes models such as fine-tuned SD and compressed SD trained using pruning and distillation techniques, among others. Below, we will introduce several downstream image generation applications based on the SD or SDXL foundation model.
Fine-tuned Checkpoint. We use two fine-tuned SDXL-based model checkpoints: “xxmix9realistic” and “samaritan-3d-cartoon”. The xxmix9realistic model is trained on person, photorealistic, sexy, female, woman, girls, and portraits using realistic style data. The samaritan-3d-cartoon model, on the other hand, is trained on anime, Disney, 3D, Pixar, niji, midjourney style.
@@
ControlNet. “controlnet-canny” is trained on SDXL with canny conditioning.
Inpainting. SDXL Inpainting 0.1 is a latent T2I diffusion model capable of generating photo-realistic images given any text input, with the extra capability of Inpainting the pictures by using a mask.
LCM. Latent Consistency Models [19] have achieved impressive performance in accelerating T2I generative tasks, producing high-quality images with minimal inference steps. “lcm-lora-sdxl” is s universal acceleration module based on SDXL.
SSD. The Segmind Stable Diffusion Model is a distilled 50% smaller version of the SDXL, offering a 60% speedup while maintaining high-quality T2I generation capabilities.
All the above-mentioned pretrained checkpoints will be detailed in Appendix 6. Through Figure 3, we have proven that the PEA module can be used as a plugin between the Chinese CLIP text encoder and the UNet, achieving language transfer ability from English to Chinese without any other process. For more detailed image generation, please refer to our Appendix 10.

### v2

@@ 4.1 Data Preparation
We download the Laion2B-multi [31] dataset along with English translations for image descriptions. Based on the open-source CLIP models available for different languages: Chinese CLIP, Russian CLIP, Japanese CLIP, Korean CLIP, and Italian CLIP, we determine to conduct experiments on the above five languages. For Chinese, we also collect WuKong [7] data with Chinese cultural image-text pairs and Chinese-English parallel corpus translations of Laion2B-en. For more data, model, and code details, please refer to the Appendix.
@@ 4.2 Implementation Details and Evaluation
Implementation details. We conduct all experiments on Stable Diffusion (SD)- [27],and Stable Diffusion XL (SDXL)- [23] like backbones. Take SDXL base model as an example. PEA-Diffusion consists of VAE, UNet, CLIP text encoder, and the adapter module, comprising 3 billion parameters, out of which a mere 6 million parameters ($\approx$ 0.2%) are trainable. We set $\lambda_{FL}$ 0.1 and $\lambda_{L}$ 1 respectively based on the loss performance. The entire model is trained on 64 A100 GPUs for 70000 steps with a batch size of 5 per GPU.
@@
Evaluation metrics. For the general evaluation data MG, the widely used FID metric can only measure the feature distance between generated images and specific images. It often fails to provide accurate assessment for high-quality generated images, as emphasized and experimentally validated by SDXL [23] and other studies. Therefore, in order to evaluate the quality of generated images more accurately, we introduce two human preference models as our evaluation metrics: ImageReward [33] and PickScore [15]. Additionally, we continue to utilize OpenCLIP ViT-bigG to measure the similarity between images and text. As for the language-specific evaluation data MC, constructed for each language, to ensure a fair comparison of content relevance, we only consider the CLIP model trained in the specific language to calculate the correlation between generated images and their corresponding text. Since the general CLIP model can only process non-English data after translation into English, this process can introduce significant performance degradation. Furthermore, our paper’s focus is precisely on addressing the errors caused by translating non-English prompts into English to generate images. Therefore, other metrics that involve translation processes, such as the two models based on human preference mentioned above, will no longer be valid for evaluation.
@@
  • Fine-tuning with LoRA [12] and $W_{K},W_{V}$ (key and value projection matrices): As low-cost experiments for comparison, we also fine-tuning the model to transfer language capabilities.
@@ 4.3 Experimental Results
Generation analysis. We validate our approach based on two different T2I base models. Firstly, to ensure a fair comparison with GlueGen and AltDiffusion, we train PEA-Diffusion based on the SD1.5 model. Secondly, to validate our low-cost experiment approach, we train a larger model based on SDXL. The compared methods are LoRA fine-tuning with rank 64 and $W_{K},W_{V}$ fine-tuning, both of which are parameter-efficient fine-tuning strategies. For the second approach, as proposed in [16], when updating the mapping from given text to image distribution, only updating $W_{K}^{(i)},W_{V}^{(i)}$ in each cross-attention block $i$ is sufficient since text features are the only input to the key and value projection matrix. Furthermore, to validate the wide application of the proposed approach with different text encoders, we train our model with a multi-language T5 encoder and a multi-language CLIP encoder. This type of training only requires English text data, which makes it even simpler and more efficient. Fig. 2 illustrates the comparisons between translation and our methods with different training strategies. It can be observed that our PEA-Diffusion can achieve Chinese culturally related image generation, but translation totally fails in these cases.
Table 1: CLIPScore with different specific languages. For Chinese, we evaluated two general evaluation metrics with MG data. $\dagger$ Indicates open source models.
@@
Altdiffusion m9 [36$\dagger$ 0.2100 -0.6966 0.1202 0.3715 0.3177 0.1906 0.2461
Altdiffusion m18 [36$\dagger$ 0.2172 0.3316 0.1648 0.4730 0.3502 0.1917 0.3008
GlueGen [24$\dagger$ 0.1700 -0.5187 0.1224 0.4184 0.3065 0.1769 0.2506
@@
As shown in the comprehensive results of Table 1, when compared with the translation method, our PEA-Diffusion outperforms it for almost all languages (except Japanese) in terms of CLIPScore for MC datasets. It is obvious that the translation method cannot capture the essence of different languages’ cultures, leading to a semantic gap between the original prompts and the translated English prompts. However, our approach can use additional culturally related image-text pairs to train the adapter, which can align the language understanding ability of the text encoder with the generation ability of the UNet. Even though the UNet is frozen during the whole training process, we argue that our PEA-Diffusion can still stimulate the potential generation ability for cultural-related images of the original UNet. Since SD1.5 and SDXL are all trained on billions of image-text pairs that contain wide image domains, our approach can successfully awaken their generative abilities in different cultural contexts. In terms of ImageReward and PickScore for the MG dataset, we must admit that PEA-Diffusion is slightly inferior to translation methods. This is because using other text encoders (e.g., Chinese CLIP and Multilingual CLIP) cannot ensure complete alignment between the text embedding and the UNet feature spaces.
@@
Even though Altdiffusion retrains the multilingual CLIP text encoder and UNet on billions of image-text pairs, it still performs worse than our approach in terms of almost all metrics for all languages. GlueGen is the most related work on our concerned topic. The statistics presented in Table 1 reveal that both checkpoints provided by the authors or re-implemented by ourselves are significantly inferior to our PEA-Diffusion, especially for the ImageReward and PickScore metrics, resulting in very low generation quality since only aligning distributions of text embedding for different languages is very hard, even with its sophisticatedly designed GlenNet. In addition, it cannot generate culturally related images. Therefore, it’s worth mentioning that our core contribution is that we align the model’s generation ability through UNet distillation only by a MLP-like adapter between different languages. We don’t forcibly align the text embedding spaces of different languages.
Additionally, we undertake various parameter-efficient fine-tuning strategies for comparison. The fine-tuning techniques, including LoRA and $W_{k}$ and $W_{v}$, show subpar performance based on ImageReward, PickScore, and CLIPScore metrics. This is attributed to the fact that these fine-tuning methods depend significantly on high-quality image-text datasets, whereas our PEA-Diffusion requires merely some parallel multilingual texts and images and lacks stringent quality requirements. Further experiments can be found in the Appendix.
@@
Analysis on different languages. The data in Table 1 indicate that the CLIPScores for Russian, Italian, and Korean either match or exceed those of other baseline methods by a notable margin. For Japanese, however, the CLIPScore is slightly lower compared to translation and AltDiffusion. We ascribe this to the CLIP model’s limited encoding ability for Japanese, which fails to align with the SD model via the adapter layer. Additionally, with only 1.6M training samples for Korean, our method still achieves effective language transfer, underscoring that large-scale training datasets are not essential. Further visualization examples of culturally-related prompts can be found in the Appendix.
Figure 3: Training parameters(M) and cost comparison for different methods.
@@
Analysis of training cost. Our PEA-Diffusion freezes the CLIP text encoder, the UNet, and VAE during the whole training process, only the MLP mapping adapter is trained, accounting for merely 6M training parameters, which is 0.2% of total parameters for SDXL-based models. We employed 64 A100 GPUs to train 70000 steps, resulting in 1600 GPU hours. As a comparison, AltDiffusion [36] trained the model for three stages with nearly 30000 GPU hours, which is $\approx 20$ times our training cost. In addition, our method achieves higher performance in terms of almost all metrics and the five experimental languages. We also re-implemented and trained GlueGen [24] with the same dataset constructed by ourselves. It indeed achieves very low training costs with only 1 A100 GPU for 24 hours, but results in very poor performance no matter for image quality (-0.5 ImageReward score) or language transferring ability. Finally, we also conduct other parameter-efficient fine-tuning strategies (e.g., LoRA and tuning only $W_{k}$ and $W_{v}$ matrices of attention layers) and achieve much lower performance and higher training costs, which demonstrated the effectiveness and efficiency of our PEA-Diffusion. Figure 3 only shows the relationship between training parameters and GPU hours of different methods. For more information on model fitting and training data, as well as adapter module parameter variations, please refer to Appendix.
4.4 Ablation Studies
Since Chinese data has certain advantages in terms of data quality and quantity, to eliminate errors caused by data, our ablation experiments are all based on Chinese data. The experiments include four parts: First, we conduct a baseline experiment where we only perform standard model fine-tuning without incorporating any KD strategies. Secondly, inspired by GlueGen [24], we employ text encoder KD to explore the impact of distillation at different positions in the SD framework. Next, we perform distillation based on the UNet model, including feature distillation between different UNet layers and logits distillation for the final output. Finally, we combine all of the distillation positions described above.
@@ 4.4 Ablation Studies
Fine-tuning -1.3154 0.1854 0.3278 0.1901
Fine-tuning w/ TKD -0.9875 0.1997 0.3506 0.2287
Fine-tuning w/ ULKD -1.0447 0.1957 0.3331 0.1968
Fine-tuning w/ UFKD 0.9556 0.2174 0.4085 0.2557
Fine-tuning w/ UTKD 0.8490 0.2113 0.4011 0.2461
Ours(Fine-tuning w/ UKD) 1.0245 0.2244 0.4224 0.2610
Effectiveness of KD. First, KD is crucial for guiding models onto the right learning trajectory for rapid convergence and is vital for the PEA-Diffusion method. As indicated in the first row of Table 2, the standalone fine-tuning method yields notably inferior performance in both image quality (ImageReward and PickScore) and image-text alignment (CLIPScores).
@@
Effectiveness of UNet Feature KD (UFKD). Our KD method consists of two parts: the first part involves distilling the logits results generated by the UNet (ULKD), we then introduce a distillation of intermediate layer features in the UNet, which serves as the primary knowledge guidance mechanism. As shown in the third and fourth lines of Table 2, we can observe that logit distillation can only improve marginally when compared with direct fine-tuning whereas feature distillation boosts the performance remarkably. We attribute this phenomenon to the generation ability of the large amount of UNet parameters, so that intermediate feature supervision is essential for knowledge transfer.
@@
Effectiveness of UNet KD. Furthermore, different distillation positions have different effects. Our initial focus is on aligning the feature spaces of different text encoders, so a natural idea is to distill only the feature information of the teacher model’s text encoder (TKD). However, as the results show in the second line of Table 2, although the overall performance exceeds the pure fine-tuning method, it is still far from our best performing approach. This illustrates that simply aligning the text encoder is insufficient and ineffective. This statistic also matches the experimental results of GlueGen in Table 1: simply aligning text embedding spaces cannot obtain satisfactory image generation results. Therefore, this further proves the necessity of distilling the features of the UNet itself.
@@
Finally, we perform KD on both the UNet and text encoder (UTKD). However, the fifth line of Table 2 shows that its effect is worse than performing only UNet distillation. We suspect that the addition of text encoder KD actually weakens the learning of the SD core UNet structure. The main knowledge ability for generating images is mainly influenced by the UNet model, which is responsible for the adapter alignment feature in PEA-Diffusion. More ablation experiments can refer to Appendix.
4.5 Exploring the Domain Adaptation of PEA-Diffusion
To demonstrate the domain adaptation of our model with minimal data guidance, we first constructed a dataset of 20,000 parallel image and text pairs in the domain of human. We then trained the models based on the models fine-tuned $W_{K}^{(i)},W_{V}^{(i)}$ mentioned in section 4.3 with UKD, and evaluated on the MG and MC dataset. The results on Chinese are shown Table 3.
Table 3: Domain adaptation with 20,000 data in Chinese specific language.
Evaluation ImageReward PickScore CLIPScore_en CLIPScore_zh
$W_{k}$ and $W_{v}$ 0.1413 0.1621 0.3674 0.2485
Ours(Fine-tuning w/ UKD) 0.8368 0.2145 0.4098 0.2501
We observed significant improvements in the three metrics on the MG test data. This suggests that PEA can be guided by a small amount of data to perform out-of-domain knowledge transfer.
@@ 4.6 Adaptability Experiment of English Downstream Model
Figure 4: Downstream task experiments of our proposed PEA. By only adding the trained adapter between the Chinese CLIP text encoder and the UNet, it can achieve language transfer for fine-tuning, LoRA, ControlNet, Inpainting, LCM and compressed models.
In this section, we would like to emphasize the plug-and-play feature of our PEA module. Currently, a wide range of community resources, such as fine-tuned models, LoRA, Dreambooth, ControlNet, compressed SD models, and accelerated sampling SD models, are derived from the English SD models, so that the English community is much more mature than other languages. Therefore, one key challenge we must overcome is bridging the compatibility gap between the new model and the English SD community after language transfer. Since our approach uses feature-level and logit-level KD to train only the PEA module while keeping the core UNet parameters unchanged, it naturally maintains compatibility with the English SD community and can seamlessly integrate with various English SD community plugins. Interestingly, the PEA module remains plug-and-play for models that have undergone changes in base SD UNet parameters in contrast to PEA-Diffusion but whose overall structure remains similar. This includes models such as fine-tuned SD and compressed SD trained using pruning and distillation techniques, among others. In the following, we will introduce several downstream image generation applications based on the SD or SDXL foundation model.
Fine-tuned checkpoint. We use two fine-tuned SDXL-based model checkpoints: “xxmix9realistic” and “samaritan-3d-cartoon”. The “xxmix9realistic” model is trained on person, photorealistic, sexy, female, woman, girls, etc. The “samaritan-3d-cartoon” model is trained on anime, Disney, 3D, Pixar, etc.
@@
ControlNet. “controlnet-canny” is trained on SDXL under canny condition.
Inpainting. SDXL Inpainting 0.1 is a latent T2I diffusion model capable of generating photorealistic images given any text input, with the extra capability of Inpainting the pictures by using a mask.
LCM. Latent Consistency Models [20] have achieved impressive performance in accelerating T2I generative tasks, producing high-quality images with minimal inference steps. “lcm-lora-sdxl” is an acceleration module based on SDXL.
SSD. The Segmind Stable Diffusion Model is a condensed version of the SDXL that is 50% smaller in size, delivering a 60% increase in speed while preserving high-quality text-to-image (T2I) generation capabilities.
In Figure 4, we have shown that the PEA module can be used as a plugin between the Chinese CLIP text encoder and the UNet, achieving the ability to transfer language from English to Chinese directly. For more details of the above-mentioned checkpoints and generated images, please refer to our Appendix.